If the momentum of an object is tripled, its kinetic energy will change by what factor?
Answer #1

\(\mathrm{p}=\mathrm{mv}\)

\(3 p=3 m v\)

\(\mathrm{K}=\frac{1}{2} m v^{2}\)

the mass is the same then \(v\) is increasing 3 times

\(\mathrm{K}=\frac{1}{2} m(3 v)^{2}\)

\(\mathrm{K}=\frac{1}{2} m(9) v^{2}\)

then 9 times